OpenStudy (anonymous):

Which of the following equations has only one solution? x^2 + 4x + 4 = 0 x^2 + x = 0 x^2 - 1 = 0

OpenStudy (jziggy):

If you're allowed to use graphing calculators, the simplest method is to graph each of the equations and see which touch the x-axis at one point only. However, as that probably won't be allowed on tests, you can also factor the equations. 1. (x+2)(x+2) = 0 x = -2 2. x(x+1) = 0 So, x = 0 and x = -1 3. x^2 = 1 So, x = -1 and x = 1